# On-Device LLMs for Home Assistant: Dual Role in Intent Detection and Response Generation

### v1

@@ Abstract
Rune Birkmose*     Nathan Mørkeberg Reece*     Esben Hofstedt Norvin* Johannes Bjerva     Mike Zhang Aalborg University, Denmark jjz@cs.aau.dk
@@ 3 Methodology
Table 1: Aggregated Train/Test Splits. For the baseline, 20% of the original training set was used as test data (after removing multi-intent samples). The LLM used the full synthetic data; 2,435 remain as test.
@@ 3.1 Data and Pre-processing
A full example can be found in Figure 1 (Appendix A). We stratify the dataset, maintaining the inherent imbalance (some device types and services appear more frequently). We split into training and test sets as shown in Table 1. The final training set for the LLM includes $\sim$33k examples, and we set aside 2,435 synthetic samples for evaluation. Note that for the baseline SVC, a different test partition was selected (5,819 samples) to exclude multi-intent examples. A more detailed distribution of the data can be found in Table 6 (Appendix B).
@@
SVC (Baseline) 84.0%
Qwen2.5-0.5B (16-bit) 98.8% 0.84
Qwen2.5-0.5B (8-bit) 98.4% 0.79
Qwen2.5-0.5B (4-bit) 81.7% 0.88
Qwen2.5-1.5B (16-bit) 96.9% 0.84
Qwen2.5-1.5B (8-bit) 96.5% 0.83
Qwen2.5-1.5B (4-bit) 90.7% 0.82
@@ 3.2 Models
Baseline Classifier.
We train a Linear SVC from Scikit-Learn (Pedregosa et al., 2011) on TF-IDF features of the user prompt. The classifier outputs a concatenated device-service pair, which is then wrapped in JSON. It has no generative capability, so user-facing text is templated.
@@
We train using a chat-style format with user–assistant pairs. We primarily use the Qwen2.5-0.5B-Instruct model and the Qwen2.5-1.5B-Instruct (Yang et al., 2024). We fine-tune both models for one epoch with a batch size of 4, using the AdamW optimizer at a learning rate of $2\times 10^{-5}$ with a cosine scheduler. The maximum sequence length is set to 2,048 tokens. We use the HuggingFace Transformers library (Wolf et al., 2020) for training on NVIDIA L4 (24 GB) Tensor Core GPUs.
@@
After fine-tuning and having the original 16-bit (BF16) model, we produce two quantized versions of each model: NF8 and NF4 Dettmers et al. (2024), using bitsandbytes. [^3] See https://github.com/bitsandbytes-foundation/bitsandbytes. We also use double quantization. This allows us to compare accuracy, generative quality, and inference speed under varying memory constraints.
3.3 Evaluation
Slot-Intent Detection Accuracy.
Text Generation Quality.
For the natural language responses, we compare each generated response to the reference using BERTScore (Zhang et al., 2020).
Inference Environment.
@@ 4.1 Slot and Intent Detection
Table 2 shows the SID performance of both the 0.5B and 1.5B LLMs under various quantization levels, alongside the baseline SVC. For the 0.5B model, the 16-bit and 8-bit variants reach near-perfect accuracy ($\sim 99\%$). The 4-bit version drops to 81.7%, which is marginally below the SVC baseline (84.0%).
@@ 5 Discussion
Despite near-perfect performance on the synthetic test set, Table 4 and 5 reveal a drop to 80–86% accuracy in real or out-of-domain queries. This discrepancy likely stems from the difficulty of handling spontaneous human phrasing, missing location or device details, and genuinely novel device types. Still, the results match or surpass the traditional SVC baseline.
@@ Appendix B Data Distribution Detailed
Table 6: Detailed Class Distribution. Total Dataset vs. LLM Test Subset

### v2

@@ Abstract
Rune Birkmose*     Nathan Mørkeberg Reece*     Esben Hofstedt Norvin* Johannes Bjerva     Mike Zhang Aalborg University, Denmark *{rbirkm20, nreece20, enorvi20}@student.aau.dk   jjz@cs.aau.dk
@@ 3 Methodology
Table 1: Aggregated Train/Test Splits. For the classification baseline, 20% of the original training set was used as test data (after removing multi-intent samples). The LLM used the full synthetic data; 2,435 remain as test.
@@ 3.1 Data and Pre-processing
A full example can be found in Figure 1 (Appendix A). We stratify the dataset, maintaining the inherent imbalance (some device types and services appear more frequently). There are 38 service labels and 858 device labels. We split into training and test sets as shown in Table 1. The final training set for the LLM includes $\sim$33k examples, and we set aside 2,435 synthetic samples for evaluation. Note that for the classification-based baselines, we split up the train and test set to separately predict service and device instead of as one prediction, ending up with double the test data (5,843 samples; excluding multi-intent examples). The input consists of only the user message and leave the system message out. A more detailed distribution of the data can be found in Table 6 (Appendix B).
@@
Baselines
SVC Classifier
      Service 76.6
      Device 45.4
DistilBERT
      Service 98.8
      Device 47.9
Qwen2.5-0.5B (16-bit) 98.8 0.84
Qwen2.5-0.5B (8-bit) 98.4 0.79
Qwen2.5-0.5B (4-bit) 81.7 0.88
Qwen2.5-1.5B (16-bit) 96.9 0.84
Qwen2.5-1.5B (8-bit) 96.5 0.83
Qwen2.5-1.5B (4-bit) 90.7 0.82
@@ 3.2 Models
Baseline Classifiers.
We train a Linear SVC from Scikit-Learn (Pedregosa et al., 2011) on TF-IDF features of the user prompt. The classifier outputs a concatenated device-service pair, which is then wrapped in JSON. Additionally, we fine-tune DistilBERT Sanh et al. (2019) for classification. We use the transformers library (Wolf et al., 2020) for fine-tuning. We train for 1 epoch using a learning rate of 3×$\times 10^{-4}$ with the AdamW optimizer, and a batch size of 64 on a NVIDIA A10 (24GB) GPUs. Both models have no generative capability, so user-facing text is templated.
@@
We train using a chat-style format with user–assistant pairs. We primarily use the Qwen2.5-0.5B-Instruct model and the Qwen2.5-1.5B-Instruct (Yang et al., 2024). We fine-tune both models for one epoch with a batch size of 4, using the AdamW optimizer at a learning rate of $2\times 10^{-5}$ with a cosine scheduler. The maximum sequence length is set to 2,048 tokens. We use the HuggingFace Transformers library (Wolf et al., 2020) for training on NVIDIA L4 (24 GB) GPUs.
@@
After fine-tuning and having the original 16-bit model, we produce two quantized versions of each model: NF8 and NF4 Dettmers et al. (2024), using bitsandbytes. [^3] See https://github.com/bitsandbytes-foundation/bitsandbytes. We also use double quantization. This allows us to compare accuracy, generative quality, and inference speed under varying memory constraints.
@@
3.3 Evaluation
Slot-Intent Detection Accuracy.
For the classification task, instead, we separately predict service and device using the same classification model and take the average accuracy.
Text Generation Quality.
For the natural language responses using the LLMs, we compare each generated response to the reference using BERTScore (Zhang et al., 2020).
Inference Environment.
Model Accuracy BERTScore
Qwen2.5-0.5B 80.0 0.76
Qwen2.5-1.5B 86.7 0.74
@@ 4.1 Slot and Intent Detection
Table 2 shows the SID performance of both the 0.5B and 1.5B LLMs under various quantization levels, alongside the baseline SVC and DistilBERT. For the 0.5B model, the 16-bit and 8-bit variants reach near-perfect accuracy ($\sim 99\%$). The 4-bit version drops to 81.7%, which is still better than the the SVC baseline (average 61.0% accuracy) and DistilBERT baseline (average 73.4% accuracy).
@@ 5 Discussion
Despite near-perfect performance on the synthetic test set, Table 4 and 5 reveal a drop to 80–86% accuracy in real or out-of-domain queries. This discrepancy likely stems from the difficulty of handling spontaneous human phrasing, missing location or device details, and genuinely novel device types. Still, the results surpass the SVC and DistilBERT baseline.
@@ Appendix B Data Distribution Detailed
Table 6: Detailed Class Distribution Service. Total Dataset vs. LLM Test Subset